# Fully Kinetic Simulations of Proton-Beam-Driven Instabilities from Parker Solar Probe Observations

### v1

@@ 3 Results
Figure 1: Results for the linear Vlasov theory, corresponding to our initial setup, in the left column, compared with the simulation results in the right column. Panel (a) to (f) show the following: the real frequency $\omega_{\mathrm{r}}/\omega_{\mathrm{p}}$ (a); the imaginary frequency $\gamma$ (c); the polarization of the wave (e). The right column shows the following: in panel (b), the power spectrum measured for a cut at $y=L_{y}/2$ is plotted against the frequency at different positions along the $x$-axis; panel (d) shows the power spectrum at $t/\Omega_{\mathrm{p}}^{-1}\approx 218$, with some of the instability growth rates contours of panel (c) overlaid, with the black dot denoting the fastest-growing mode; panel (f) is the hodogram at $t/\Omega_{\mathrm{p}}^{-1}\approx 218$, where the red triangle shows the value of the magnetic fluctuation corresponding to the point $x=L_{x}$, see the text for details on how this is computed.
@@
The results from our numerical simulation are compared with the Vlasov linear theory in Figure 1. We solved the linear dispersion relation using the DIS-K solver [^1] The code is publicly available and can be found at https://github.com/ralopezh/dis-k. for the setup described above (López et al., 2021; López, 2023). Specifically, the spectrum and polarisation of the waves developed by the system are analyzed. The real part $\omega_{\mathrm{r}}/\omega_{\mathrm{p}}$ of the complex frequency , as obtained from linear theory, is illustrated in Figure 1(a) against $k_{\parallel}$ and $k_{\perp}$. In Figure 1(b) we plot the power spectrum of the magnetic fluctuations $\delta\boldsymbol{B}/B_{0}=(\boldsymbol{B}-\boldsymbol{B}_{0})/B_{0}$ measured along the parallel direction fixing $y=L_{y}/2$, denoted as $\|\mathcal{F}\left\{\delta\boldsymbol{B}/B_{0}(y=L_{y}/2)\right\}\|^{2}$. Here, $\mathcal{F}$ represent the Fast Fourier Transform (FFT) in time, and the colorbar indicates the variation in the spatial dimension $x/L_{x}$. The peak of maximum intensity occurs at $\omega_{\mathrm{r}}/\omega_{\mathrm{p}}=0.00151$ at all $x/L_{x}$ as pointed out in panel (b) by the black dashed line, corresponding to $k_{\parallel}/d_{\mathrm{p}}^{-1}\approx 0.501$ in panel (a). According to Vlasov linear theory, the imaginary part of the complex frequency, $\gamma$, represents the growth rate of the unstable modes. Figure 1(c) plots the power spectrum of $\gamma/\omega_{\mathrm{p}}$ in the $k_{\parallel}$$k_{\perp}$ plane. In Figure 1(d), we present the analog of panel (c) for the simulated data. We display the power spectrum of $\delta\boldsymbol{B}/B_{0}$ at $t/\Omega_{\mathrm{p}}^{-1}\approx 218$. Similarly to panel (b), $\mathcal{F}\left\{\delta\boldsymbol{B}/B_{0}\right\}$ represents the 2D FFT in the spatial domain of $\delta\boldsymbol{B}/B_{0}$. Black isocontours of $\gamma/\omega_{\mathrm{p}}$ values from linear theory in panel (c), have been superimposed to highlight the good agreement between analytic and simulation results. This specific time has been chosen because it corresponds to the moment when the magnetic energy reaches its maximum, see Figure 4(a). The dominant mode (i.e. fastest-growing mode), indicated by the black dot, propagates mostly along the parallel direction ($x$-axis) with $k_{\parallel}/d_{\mathrm{p}}^{-1}\approx 0.501$ and $k_{\perp}/d_{\mathrm{p}}^{-1}\approx 0.0245$, having a magnitude of $\gamma_{\mathrm{m}}/\Omega_{\mathrm{p}}\approx 0.042$. In Figure 1(e), we plot the polarisation $P$ of the electric field of the wave as calculated from linear theory. If $P>0$, then the wave is right-handed (RH) elliptically polarised, whereas if $P<0$, it is left-handed (LH). In this case, the polarisation is RH, but it is very close to unity, making the expected hodogram almost circular. The hodogram in Figure 1(f) provides information about the polarisation of the wave for our simulation. By the reported time frame of $t/\Omega_{\mathrm{p}^{-1}}\approx 218$ all unstable modes have completed their development. A wave polarisation state can be split into two linearly polarized orthogonal components with respect to the guiding magnetic field (Narita, 2017; Born & Wolf, 2019). In our case, the two components are measured as the values, along the $x$-axis at $y=L_{y}/2$, of $\delta B_{y}/B_{0}$ and $\delta B_{z}/B_{0}$, representing the magnetic fluctuation in the transverse directions at $t/\Omega_{\mathrm{p}}^{-1}\approx 218$. The red triangle highlights the value of the magnetic fluctuation at $x/{d_{\mathrm{p}}}=L_{x}$ and $y/{d_{\mathrm{p}}}=L_{y}/2$. From this point, a counterclockwise rotation corresponds to an RH circular polarization state, while a clockwise rotation refers to an LH circular polarization state. In Figure 1(f) we have an RH circularly polarized wave, which once again aligns with the linear theory shown in Figure 1(e) and discussed in Klein et al. (2021).
@@
Figure 2: Magnetic field fluctuations resulting from the PIC simulation. Panel (a) displays the spectral power of magnetic fluctuations for different wavenumbers (i.e., different modes), corresponding to those within the region delimited by the inner contour in Figure 1(d) with $k_{\parallel}/d_{\mathrm{p}}^{-1}\in[0.3,0.6]$. The calculated fastest-growing mode, represented by the violet line and highlighted with its corresponding wavenumber, is compared with the expected maximum growth rate from the linear theory, indicated by the dashed black line. In Panel (b), a spatio-temporal analysis is presented, focusing on the cut along $y=L_{y}/2$ of the magnetic fluctuation stacked in time. The black line in panel (b) has a slope equal to the phase velocity of the wave calculated from linear theory.
In Figure 2(a), the spectral power of magnetic fluctuations has been plotted for for a number of unstable modes within the range of parallel wavenumbers $k_{\parallel}/d_{\mathrm{p}}^{-1}\in[0.3,0.6]$ Figure 1(d). Our main focus is on the mode with the steepest slope, representing the fastest-growing one. This mode possesses wavenumber values $k_{\parallel}/d_{\mathrm{p}}^{-1}=0.501$ and $k_{\perp}/d_{\mathrm{p}}^{-1}=0.0245$ and along with the growth rate $\gamma_{\mathrm{m}}/\Omega_{\mathrm{p}}=0.042$, as determined by linear theory. The maximum theoretical growth rate $\gamma_{\mathrm{m}}$ from the linear analysis of Figure 1(c) is drawn in dashed line as comparison. The fastest growing mode is exponentially growing during the linear phase of the instability, occurring from time $t/\Omega_{\mathrm{p}}^{-1}\approx 116$ to $t/\Omega_{\mathrm{p}}^{-1}\approx 189$, closely matching the predicted growth rate according to Vlasov linear theory. From the analysis in Figure 1(b)-(c), we can calculate the phase velocity of the wave corresponding to the most-unstable mode, denoted as $v_{\mathrm{ph}}=\omega_{\mathrm{r}}/k_{\parallel}\approx 0.00306$. In Figure 2(b), a spatial and temporal analysis of the magnetic fluctuations $\delta\boldsymbol{B}/B_{0}(y=L_{y})$ is presented. This plot was obtained stacking in time the value of $\delta\boldsymbol{B}/B_{0}(y=L_{y})$. The line we overlaid, with equation $x=v_{\mathrm{ph}}t+b$, where $b/d_{\mathrm{p}}\approx-130.6$ is the offset, demonstrates that the perturbation in the magnetic field is propagating at the speed predicted by the linear analysis.
@@
Figure 3: Fluctuations in the proton density $\rho_{\mathrm{p}}$ in panel (a), with $\delta\rho_{\mathrm{p}}=\rho_{\mathrm{p}}(t)-\rho_{0}$, and fluctuations in the magnetic field components $B_{x},B_{y},B_{z}$, in panel (b)-(d), at the moment of the instability saturation $t/\Omega_{\mathrm{p}}^{-1}\approx 218$. The amplitude of $\rho_{\mathrm{p}}$ and $B_{x}$ fluctuations, panels (a) and (b), have been multiplied by a factor 10 for visualisation purposes.
@@
The spatial distribution of fluctuations in proton mass density $\rho_{\mathrm{p}}$ and magnetic field components is displayed in Figure 3(a)–(d). They are presented at time $t/\Omega_{\mathrm{p}}^{-1}\approx 218$, corresponding to the moment of instability saturation after the end of the linear phase, to examine the waves when they are fully developed. The perturbation in proton density, as visible in panel (a), indicates that the instability develops as a compressive wave. Additionally, there is a positive correlation between the fluctuation in proton density and the magnetic field fluctuation in the $x$-direction, as denoted by both the pattern and intensity shown in Figure 3(a)-(b). The magnetic fluctuation in the $y$ and $z$ directions, representing the transverse directions, is more intense, indicating that it carries more energy than that in the parallel direction. We observe that these wavepackets are predominantly parallel-propagating, and thus, $\theta\doteq k_{\perp}/k_{\parallel}\approx 0.05$, with $\theta$ representing the angle between the wavevector components $k_{\parallel}$ and $k_{\perp}$ as calculated from linear theory. Therefore, we can infer that the developed wave is a fast-magnetosonic/whistler (FM/W) wave.
@@
Figure 4: Evolution of different types of energy and proton temperature ratio. Panels (a)-(c) illustrate the time evolution of different energy components, categorised into three distinct phases named I, II, and III, corresponding to the stationary, linear, and nonlinear phases, respectively. The gray region identifies phase II, with the dashed-dotted line indicating the inflection point of the curves, where the conversion of kinetic energy into magnetic energy peaks at $t_{\mathrm{peak}}/\Omega_{\mathrm{p}}^{-1}\approx 170$. Panel (a) displays the changes in magnetic and total kinetic energy of electrons and protons, normalised by the initial total energy of the system. Panel (b) presents the thermal energies (parallel and perpendicular with respect to the guide field direction) of both the core and beam populations, along with the core and beam drift (i.e. bulk) energy. In panel (c), the temporal evolution of the temperature ratio of the total proton population, encompassing both the core and beam, is shown. The formal definition of these quantities is provided in the text.
@@
In Figure 4, we show the energy exchanges within the considered system. In Figure 4 (a)-(c) we identify three phases of evolution the system is passing through. Phase I or quasi-stationary phase ends at approximately $t_{\mathrm{beg}}/\omega_{\mathrm{p}}^{-1}\approx 126$, which coincides with the moment the instability starts growing with the fastest growing mode. Subsequently, Phase II or linear phase, in light gray, begins at the end of phase I and evolves until the instability saturates and then the conversion from magnetic to kinetic energy reaches its maximum around $t_{\mathrm{end}}/\omega_{\mathrm{p}}^{-1}\approx 251$ as illustrated in Figure 2(a). At the end of this phase, the total magnetic and kinetic energies respectively reach their maximum and minimum values. Moreover, the dash-dotted line in Figure 4 highlights the inflection point of the energies, which is the point where the conversion of kinetic energy into magnetic energy peaks at $t_{\mathrm{peak}}/\omega_{\mathrm{p}}^{-1}\approx 170$. It is followed by Phase III, named non-linear phase, which starts at $t_{\mathrm{end}}/\omega_{\mathrm{p}}^{-1}\approx 251$ and continues until the end of the simulation.

### v2

@@ 3 Results
Figure 1: Results from linear Vlasov theory corresponding to our initial setup, in the left column, compared with the simulation results in the right column. Panels (a) to (f) show the following: the real frequency $\omega_{\mathrm{r}}/\omega_{\mathrm{p}}$ (a); the imaginary frequency $\gamma$ (c); the polarization of the wave (e). The right column shows the following: in panel (b), the power spectrum measured for a cut at $y=L_{y}/2$ is plotted against the frequency at different positions along the $x$-axis, the black and red dashed lines represent the frequencies obtained from the linear theory and the numerical simulation, respectively; panel (d) shows the power spectrum at $t/\Omega_{\mathrm{p}}^{-1}\approx 218$, with some of the instability growth rates contours of panel (c) overlaid, with the black dot denoting the fastest-growing mode; panel (f) is the hodogram at $t/\Omega_{\mathrm{p}}^{-1}\approx 218$, where the red triangle shows the value of the magnetic fluctuation corresponding to the point $x=L_{x}$ (see text for details on how this is computed).
@@
The results from our numerical simulation are compared with Vlasov linear theory in Figure 1. We solved the linear dispersion relation using the DIS-K solver [^1] The code is publicly available and can be found at https://github.com/ralopezh/dis-k. for the setup described above (López et al., 2021; López, 2023). Specifically, the spectrum and polarization of the waves developed from our initial setup are analyzed. The real part $\omega_{\mathrm{r}}/\omega_{\mathrm{p}}$ of the complex frequency, as obtained from linear theory, is shown in Figure 1(a) against $k_{\parallel}$ and $k_{\perp}$. In Figure 1(b) we plot the power spectrum of magnetic fluctuations $\delta\boldsymbol{B}/B_{0}=(\boldsymbol{B}-\boldsymbol{B}_{0})/B_{0}$, measured along the parallel direction fixing $y=L_{y}/2$, denoted as $\delta\boldsymbol{B}(y=L_{y}/2)/B_{0}$. Here, $\mathcal{F}$ represent the Fast Fourier Transform (FFT) in time, and the colorbar indicates the variation in the spatial dimension $x/L_{x}$. In our simulation, the peak of maximum intensity occurs at $\omega^{\dagger}_{\mathrm{r}}/\omega_{\mathrm{p}}=0.00159$ in red, while the value calculated via quasi-linear theory is $\omega_{\mathrm{r}}/\omega_{\mathrm{p}}=0.00151$ in black, as pointed out in panel (b) by the black dashed line, corresponding to $k_{\parallel}/d_{\mathrm{p}}^{-1}\approx 0.501$ in panel (a). Comparisons between linear theory and simulation results reveal small discrepancies, as the VDF remains constant in linear theory but is deformed as it evolves during the course of the simulation.
Figure 1(c) shows the power spectrum of $\gamma/\omega_{\mathrm{p}}$ in the $k_{\parallel}$$k_{\perp}$ plane. The dominant mode (i.e. fastest-growing mode), indicated by the black dot, propagates mostly along the parallel direction ($x$-axis) with $k_{\parallel}/d_{\mathrm{p}}^{-1}\approx 0.501$ and $k_{\perp}/d_{\mathrm{p}}^{-1}\approx 0.0245$, having a magnitude of $\gamma_{\mathrm{m}}/\Omega_{\mathrm{p}}\approx 0.042$. In Figure 1(d), we present the analog of panel (c) for the simulated data. We display the power spectrum of $\delta\boldsymbol{B}/B_{0}$ at $t/\Omega_{\mathrm{p}}^{-1}\approx 218$ (end of the quasi-linear phase). Similarly to panel (b), $\mathcal{F}\left\{\delta\boldsymbol{B}/B_{0}\right\}$ represents the 2D FFT in the spatial domain of $\delta\boldsymbol{B}/B_{0}$. Black isocontours of $\gamma/\omega_{\mathrm{p}}$ obtained from linear theory, as in panel (c), have been superimposed in panel (d) to highlight the good agreement between analytic and simulation results. This specific time has been chosen because it corresponds to the moment when the magnetic energy reaches its maximum, see Figure 4(a).
In Figure 1(e), we plot the polarization $P\doteq\mathbb{R}\big{\{}i{\rm\operatorname{sign}}(\omega_{\mathrm{r}})D_{x}/D_{y}\big{\}}$ (Stix, 1962; Gary, 1993), calculated from linear theory, where $D_{x}$ and $D_{y}$ are the electric field component of the wave. If $P>0$, then the wave is right-handed (RH) elliptically polarised, whereas if $P<0$, it is left-handed (LH). In this case, the polarization is RH, but it is very close to unity, making the expected hodogram almost circular. The hodogram in Figure 1(f) provides information about the polarization of the wave for our simulation. By the reported time frame of $t/\Omega_{\mathrm{p}}^{-1}\approx 218$ all unstable modes have completed their development. A wave polarization state can be split into two linearly polarized orthogonal components with respect to the guiding magnetic field (Narita, 2017; Born & Wolf, 2019). In our case, the two components are measured as the values, along the $x$-axis at $y=L_{y}/2$, of $\delta B_{y}/B_{0}$ and $\delta B_{z}/B_{0}$, representing the magnetic fluctuation in the transverse directions at $t/\Omega_{\mathrm{p}}^{-1}\approx 218$. The red triangle highlights the value of the magnetic fluctuation at $x/{d_{\mathrm{p}}}=L_{x}$ and $y/{d_{\mathrm{p}}}=L_{y}/2$. From this point, a counterclockwise rotation corresponds to an RH circular polarization state, while a clockwise rotation implies an LH circular polarization state. In Figure 1(f) we have an RH circularly polarized wave, which once again aligns with the linear theory shown in Figure 1(e) and discussed in Klein et al. (2021).
@@
Figure 2: Magnetic field fluctuations resulting from the PIC simulation. Panel (a) displays the spectral power of magnetic fluctuations for different wavenumbers (i.e., different modes), corresponding to those within the region delimited by the innermost contour in Figure 1(d) with $k_{\parallel}/d_{\mathrm{p}}^{-1}\in[0.3,0.6]$. The calculated fastest-growing mode, represented by the violet line and highlighted with its corresponding wavenumber, is compared with the expected maximum growth rate from linear theory, indicated by the dashed black line. In Panel (b), a spatio-temporal analysis is presented, focusing on the cut along $y=L_{y}/2$ of the magnetic fluctuation stacked in time. The black line in panel (b) has a slope equal to the phase velocity of the wave calculated from linear theory.
@@
Figure 3: Fluctuations in the proton density $\rho_{\mathrm{p}}$ in panel (a), with $\delta\rho_{\mathrm{p}}=\rho_{\mathrm{p}}(t)-\rho_{0}$, and fluctuations in the magnetic field components $B_{x},B_{y},B_{z}$, in panel (b)-(d), at the moment of instability saturation $t/\Omega_{\mathrm{p}}^{-1}\approx 218$. The amplitude of $\rho_{\mathrm{p}}$ and $B_{x}$ fluctuations have been multiplied by a factor 10 for visualization purposes.
In Figure 2(a), the spectral power of magnetic fluctuations has been plotted for a number of unstable modes within the range of parallel wavenumbers $k_{\parallel}/d_{\mathrm{p}}^{-1}\in[0.3,0.6]$ considered in Figure 1(d). Our main focus is on the mode with the steepest growth, representing the fastest-growing one. This mode possesses wavenumber $k_{\parallel}/d_{\mathrm{p}}^{-1}=0.501$ and $k_{\perp}/d_{\mathrm{p}}^{-1}=0.0245$ along with the growth rate $\gamma_{\mathrm{m}}/\Omega_{\mathrm{p}}=0.042$, as determined by linear theory. The maximum theoretical growth rate $\gamma_{\mathrm{m}}$ from the linear analysis of Figure 1(c) is shown as a dashed line for comparison. The fastest growing mode is exponentially growing during the quasi-linear phase of the instability, occurring from time $t/\Omega_{\mathrm{p}}^{-1}\approx 116$ to $t/\Omega_{\mathrm{p}}^{-1}\approx 189$, closely matching the predicted growth rate according to Vlasov linear theory. From the analysis in Figure 1(b)-(c), we can calculate the phase velocity of the wave corresponding to the most-unstable mode, denoted as $v_{\mathrm{ph}}=\omega_{\mathrm{r}}/k_{\parallel}\approx 0.00306$. In Figure 2(b), is presented a spatial and temporal analysis $\delta\boldsymbol{B}/B_{0}(y=L_{y}/2)$. This plot was obtained by stacking in time the values of $\delta\boldsymbol{B}/B_{0}(y=L_{y}/2)$. The line we overlaid, with equation $x=v_{\mathrm{ph}}t+b$, where $b/d_{\mathrm{p}}\approx-130.6$ is the offset, demonstrates that the perturbation in the magnetic field is propagating at the speed predicted by the linear analysis.
@@
The spatial distribution of fluctuations in proton mass density $\rho_{\mathrm{p}}$ and magnetic field components is displayed in Figure 3(a)–(d). They are presented at time $t/\Omega_{\mathrm{p}}^{-1}\approx 218$, corresponding to the moment of instability saturation after the end of the quasi-linear phase, to examine the waves when they are fully developed. The perturbation in proton density, as visible in panel (a), indicates that the instability develops as a compressive wave. Additionally, there is a positive correlation between the fluctuation in proton density and the magnetic field fluctuation in the $x$-direction, as denoted by both the pattern and intensity shown in Figure 3(a)-(b). The magnetic fluctuation in the $y$- and $z$-directions, representing the transverse directions, is more intense, indicating that it carries more energy than that in the parallel direction. We observe that these wave packets are predominantly parallel-propagating, and thus $\theta\doteq k_{\perp}/k_{\parallel}\approx 0.05$, with $\theta$ representing the angle between the wavevector components $k_{\parallel}$ and $k_{\perp}$ as calculated from linear theory. Therefore, we can conclude that the developed wave is a fast-magnetosonic/whistler (FM/W) wave.
@@
Figure 4: Evolution of different types of energy and proton temperature anisotropy. Panels (a)-(c) illustrate the time evolution of different energy components. In panels (d)-(f), the temporal evolution of the temperature anisotropy for the proton core, proton beam, and the total proton population (encompassing both core and beam) is displayed. The three distinct phases of the instability—labeled I, II, and III—corresponding to the stationary, linear, and non-linear phases, are also depicted. The gray region identifies phase II, with the dashed-dotted line indicating the inflection point of the curves, where the conversion of kinetic energy into magnetic energy peaks at $t_{\mathrm{peak}}/\Omega_{\mathrm{p}}^{-1}\approx 170$. Panel (a) displays the changes in magnetic and total kinetic energy of electrons and protons, normalized by the initial total energy of the system. Panel (b) presents the thermal energies (parallel and perpendicular with respect to the guide field direction) of both the core and beam populations, along with the core and beam drift (i.e. bulk) energy. In panel (c), the thermal and bulk energies of the electron population are shown.
The formal definition of these quantities is provided in the text.
@@
In Figure 4, we show the energy exchanges within the considered system. In Figure 4 (a)-(f) we identify three phases of evolution the system is passing through. Phase I or quasi-stationary phase ends at approximately $t_{\mathrm{beg}}/\Omega_{\mathrm{p}}^{-1}\approx 126$, which coincides with the moment the instability starts growing with the fastest growing mode. Subsequently, Phase II or quasi-linear phase, in light gray, begins at the end of phase I and evolves until the instability saturates. Then, the conversion from magnetic to kinetic energy reaches its maximum around $t_{\mathrm{end}}/\Omega_{\mathrm{p}}^{-1}\approx 251$ as illustrated in Figure 2(a). At the end of this phase, the total magnetic and kinetic energies respectively reach their maximum and minimum values. Moreover, the dash-dotted line in Figure 4 highlights the inflection point of the energy variation, which is the point where the conversion of kinetic energy into magnetic energy peaks at $t_{\mathrm{peak}}/\Omega_{\mathrm{p}}^{-1}\approx 170$. It is followed by Phase III, named non-linear phase, which starts at $t_{\mathrm{end}}/\Omega_{\mathrm{p}}^{-1}\approx 251$ and continues until the end of the simulation.